# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v11

@@ 3.1 Definition of Information
Corollary 1 If $f$ is an enabled mapping from $X$ to $Y$, then $|X|\geq|Y|$, where $|X|$ and $|Y|$ denote the cardinalities of the sets $X$ and $Y$, respectively.
If $f(X)\trianglerighteq Y$, we also denote this as $f^{-1}(y)\trianglelefteq x$ or $x\trianglerighteq f^{-1}(y)$, to maintain consistency in the directionality of enabling. Fig.1(a) intuitively illustrates the dual nature of an enabling mapping, encompassing both its mathematical mapping and physical enabling attributes.
Figure 1: Schematic Diagram of Enabling Mapping, Noisy Information, and Information Transmission Chain. (a) The scene of penguin activity $X$ is mapped by an enabling mapping $f$ to an image $Y$ on a laptop; between $X$ and $Y$ there exists both a mathematical mapping relationship and a causal relationship. Moreover, mapping $X$ to $Y$ requires corresponding physical conditions and the expenditure of a certain amount of energy. (b) The left panel depicts the ontological state of noiseless information, while the right panel shows a noisy information state: the set $S_{o}^{-}$ is removed from the noiseless ontological state, and the previously disjoint set $S_{o}^{+}$ is superimposed, resulting in a noisy ontological state. (c) In the penguin activity information transmission chain, the carrier state of any subsequent piece of information can be regarded as the result of an enabling mapping from any prior ontological information state, and can also be reasonably interpreted by means of this ontological state and the corresponding enabling mapping.
Based on this, a more rigorous theoretical foundation for the definition of information can be established:
@@
Definition 2 (Mathematical Definition of Information) Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\mathbb{\subseteq O}\cup\mathbb{S}$, time of occurrence $T_{h}\mathbb{\subseteq T}$, ontological state set $S_{o}\left(o,T_{h}\right)$, carrier $c\mathbb{\subseteq O}$, reflection time $T_{m}\mathbb{\subseteq T}$, and carrier reflection set $S_{c}\left(c,T_{m}\right)$ are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I=\left\langle o,T_{h},S_{c},c,T_{m},S_{o}\right\rangle$. The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world (Wiener, 2019).
It should be noted that both $S_{o}\left(o,T_{h}\right)$ and $S_{c}\left(c,T_{m}\right)$ can each be represented as the interpretation of a set of WFFs within a formal system. However, the respective formal systems need not be identical, which constitutes one of the essential characteristics of information—its capacity to establish connections among all entities in the world, regardless of differences in their underlying formal systems.
(Shannon, 1948) is regarded as a foundational work in information theory. The paper proposed that ” $H=-\sum{p_{i}\log p_{i}}$ plays a central role as the measure of information in information theory.” The discussion in this paper includes both the source and channel of information. The source $Is$ is the origin of the information, and it generates events $e_{1}$, $e_{2},$…, $e_{n}$ with probabilities $p_{1}$, $p_{2},$…, $p_{n}$, respectively. The channel $Ic$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $Is$, the channel $Ic$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$ (where $1\leq i\leq n,\ \tau>0$). We define the ternary predicate ${IsEvent}^{3}(Is,T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $Is$ at time $T_{h}$, and the ternary predicate ${IcCode}^{3}(Ic,T_{m},c_{i})$ to represent that the channel $Ic$ received and transmitted code $c_{i}$ in the time interval $T_{m}$. Then,
$S_{Is}(Is,T_{h})={IsEvent}^{3}(Is,T_{h},e_{i})$ (3.3)
and
$S_{Ic}(Ic,(T_{h},T_{h}+\tau))={IcCode}^{3}(Ic,(T_{h},T_{h}+\tau),c_{i})$ (3.4)
are the states of the source $Is$ at time $T_{h}$ and the state of the channel $Ic$ in the time interval $(T_{h},T_{h}+\tau)$, respectively (where $(1\leq i\leq n)$). Furthermore,
$I:S_{Is}(Is,T_{h})\rightrightarrows S_{Ic}(Ic,(T_{h},T_{h}+\tau))$ (3.5)
represents the enabling mapping from the source state $S_{Is}(Is,T_{h})$ to the channel state $S_{Ic}(Ic,(T_{h},T_{h}+\tau))$. Therefore, although Shannon’s information theory investigates the various probability distributions of $S_{Is}(Is,T_{h})$ and the complex implementation mechanisms of $S_{Ic}(Ic,(T_{h},T_{h}+\tau))$, it forms a rich and extensive theoretical system. The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
@@ 3.2 Noisy Information
The realization of information is often influenced by noise or other external factors. Therefore, the following concept is introduced:
Definition 3 (Noisy Information) For information $I=\left\langle o,T_{h},S_{o},c,T_{m},S_{c}\right\rangle$, let $\widetilde{S_{o}}\left(o,T_{h}\right)$ be another state of the ontology $o$ at time $T_{h}$, and $\widetilde{S_{c}}\left(c,T_{m}\right)$ be another state of the carrier $c$ at time $T_{m}$,
@@
where the state $S_{o}^{-}\left(o,T_{h}\right)$ is the loss state of $S_{o}\left(o,T_{h}\right)$, satisfying $S_{o}^{-}\left(o,T_{h}\right)\subseteq S_{o}\left(o,T_{h}\right)$ and not being enabled to map to $\widetilde{S_{c}}\left(c,T_{m}\right)$, and the state $S_{o}^{+}\left(o,T_{h}\right)$ is the superposed state of $S_{o}\left(o,T_{h}\right)$, satisfying $S_{o}\left(o,T_{h}\right){\cap S}_{o}^{+}\left(o,T_{h}\right)=\varnothing$ and being enabled to map to $\widetilde{S_{c}}\left(c,T_{m}\right)$. The enabling mapping is then defined as:
$\widetilde{I}:\widetilde{S_{o}}\left(o,T_{h}\right)\rightrightarrows\widetilde{S_{c}}\left(c,T_{m}\right)$ (3.7)
which is the noisy information of $I$, while $I$ is its own noiseless information. Thus, the noisy information $\widetilde{I}$ is also information. It has the same ontology, time of occurrence, carrier, and reflection time as $I$, but the states of the entity and carrier may overlap but are not completely the same. Fig.1(b) illustrates the structural characteristics of noisy information.
In fact, Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenario (Shannon, 1948). In this case, he assumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
$E=f(S,N)$ (3.8)
where $f$ is the distortion function of the channel to the transmitted signal. According to OIT's view, the signal $S$ here is the ontological state of information $I$, $S_{Is}\left(Is,T_{h}\right)$. The distortion function $f(S,N)$ for signal $S$ and noise $N$ is equivalent to removing a state $S_{Is}^{-}\left(Is,T_{h}\right)$ from the source state $S_{Is}\left(Is,T_{h}\right)$ and adding another state $S_{Is}^{+}\left(Is,T_{h}\right)$. The output $E$ is then the channel's output state $\widetilde{S_{Ic}}\left(Ic,(T_{h},T_{h}+\tau)\right)$, where $(T_{h},T_{h}+\tau)$ is the time interval of the channel output. Thus,
$\widetilde{I}:\widetilde{S_{Is}}\left(Is,T_{h}\right)\rightrightarrows\widetilde{S_{Ic}}\left(Ic,(T_{h},T_{h}+\tau)\right)$ (3.9)
is the noisy information of
$I:S_{Is}\left(Is,T_{h}\right)\rightrightarrows S_{Ic}\left(Ic,(T_{h},T_{h}+\tau)\right)$ (3.10)
[…]
In this way, we use the definition of noisy information to explain Shannon's model of noisy channels in (Shannon, 1948), demonstrating that OIT is applicable to the complex scenarios of information research under noisy conditions.
@@ 3.3 Information Chain
(Xu et al., 2014) discusses the transmissibility of information through composite mappings. Based on Definitions 1 and 2, the fundamental laws governing information transmission can be characterized more rigorously as follows:
Lemma 1 (Transmissibility of Information): Let $I_{0}=\left\langle o,T_{h},S_{o},c_{1},T_{m_{1}},S_{c_{1}}\right\rangle$ and $I_{1}=\left\langle c_{1},T_{m_{1}},S_{c_{1}},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ be two pieces of information. Then, $I_{2}=\left\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ is also information. We refer to this as the information $I_{0}$ being transmitted from carrier $c_{1}$ to carrier $c_{2}$ via $c_{1}$.
Proof: According to Definitions 1 and 2, as well as the given conditions, for any $S_{o}(x,t)\in S_{o}\left(o,T_{h}\right)$, there exists $S_{c_{1}}\left(x_{1},t_{1}\right)\in S_{c_{1}}\left(c_{1},T_{m_{1}}\right)$such that $I_{0}\left(S_{o}(x,t)\right)\trianglerighteq S_{c_{1}}\left(x_{1},t_{1}\right)$. Furthermore, there exists $S_{c_{2}}\left(x_{2},t_{2}\right)\in S_{c_{2}}\left(c_{2},T_{m_{2}}\right)$ such that $I_{1}\left(S_{c_{1}}\left(x_{1},t_{1}\right)\right)\trianglerighteq S_{c_{2}}\left(x_{2},t_{2}\right)$. Define $I_{2}\left(S_{o}(x,t)\right)=S_{c_{2}}\left(x_{2},t_{2}\right)$. Since the physical realization of $S_{c_{2}}\left(x_{2},t_{2}\right)$ is contingent upon the existence of $S_{c_{1}}\left(x_{1},t_{1}\right)$, and the physical realization of $S_{c_{1}}\left(x_{1},t_{1}\right)$ in turn depends on $S_{o}(x,t)$, by causal induction, the physical realization of $S_{c_{2}}\left(x_{2},t_{2}\right)$ ultimately depends on $S_{o}(x,t)$.
Conversely, for any $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{1}(S_{c_{1}}(x_{1},t_{1}))\trianglerighteq S_{c_{2}}(x_{2},t_{2})$; and there also exists $S_{o}(x,t)\in S_{o}(o,T_{h})$ such that $I_{0}(S_{o}(x,t))\trianglerighteq S_{c_{1}}(x_{1},t_{1})$. Thus, $I_{2}(S_{o}(x,t))={I_{1}(I}_{0}(S_{o}(x,t)))\trianglerighteq I_{1}(S_{c_{1}}\left(x_{1},t_{1}\right))\trianglerighteq S_{c_{2}}(x_{2},t_{2})$.
Therefore, $I_{2}=\left\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ is an enabling mapping from $S_{o}(o,T_{h})$ to $S_{c_{2}}(c_{2},T_{m_{2}})$, which satisfies the definition of information as given in Definition 2. In essence, this is the information $I_{0}$ transmitted from carrier $c_{1}$ to carrier $c_{2}$. The lemma is thus proved.
Corollary 2 (Information Chain) Let $\left\{I_{i}=\left\langle o_{i},T_{h_{i}},S_{o_{i}},c_{i},T_{m_{i}},S_{c_{i}}\right\rangle\middle|i=1,\cdots,n\right\}$ be a sequence of information mappings, where for any $i>1$, it holds that $o_{i}=c_{i-1}$, $T_{h_{i}}=T_{m_{i-1}}$, and $S_{o_{i}}\left(o_{i},T_{h_{i}}\right)=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$. In this case, we say that these mappings constitute an information chain. Moreover, for any 1 $\leq i\leq j\leq n$, $I_{ij}=\left\langle o_{i},T_{h_{i}},S_{o_{i}},c_{j},T_{m_{j}},S_{c_{j}}\right\rangle$ also qualifies as information according to Definition 2.
Transmission constitutes a fundamental mode of information flow and application. Communication, processing, input, and output are all concrete realizations of information transmission. In many cases, information undergoes more complex flows through chain-like transmission processes. Fig.1(c) intuitively illustrates such a process. For any pair of states within this process, the latter can be regarded as the result of the enabling mapping from the former, and this enabling mapping can be used to reasonably explain the causation of the latter state. This provides a theoretical foundation for the interpretability of the model.
@@ 4 Meta-Framework for Machine Learning Theory (MLT-MF)
The core of machine learning is the artificial intelligence model, which can be regarded as a finite automaton as well as an information system. The processes of operation, learning, testing, and application for such a model all fundamentally depend on information within the system. Therefore, by employing the concepts of information model and finite automaton state (Appendix B), the process of machine learning and its applications can be unified into five primary steps of information realization and functioning:
@@ 4.1 Construction of Initial Information
Suppose the model under study is denoted as $M$, and let $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ be the set of time points at which $M$, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order. According to the perspective of OIT, the creation of $M$ is essentially the realization of information
$Io=\left\langle M_{D},{To}_{h},S_{M_{D}},M,t_{1},S_{M}\right\rangle$ (4.1)
where the ontology $M_{D}$ of $Io$ is the designer’s conceptualization of $M$, ${To}_{h}$ is the time at which the designer conceives $M$, and $S_{M_{D}}\left(M_{D},{To}_{h}\right)$ is the initial finite automaton state that must be realized at this conceptual stage. The carrier $M$ of $Io$ is the model itself; the reflection time $t_{1}$ is when $M$ is constructed according to $S_{M_{D}}\left(M_{D},{To}_{h}\right)$; and $S_{M}\left(M,t_{1}\right)$ is the formalization of the actual physical state of $M$ upon its completion. Thus, $\sup\left({To}_{h}\right)<sup(t_{1})$.
@@
$\begin{split}{\varphi o}_{2}=&{State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)\land{Input}^{3}(M_{D},t_{1},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &\land{Learnable}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))\rightarrow\exists f_{learn}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &({Eq}^{2}(\Phi^{2}(M_{D},t_{2}),f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))))\land{State}^{3}(M_{D},t_{2},\Phi^{2}\left(M_{D},t_{2}\right))\end{split}$ (4.3)
@@
$\varphi_{O_{3}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},To_{h}))\wedge Input^{3}(M_{D},t_{1},S_{M},(M_{I},t_{1}))$ (4.4)
$\wedge Processable^{2}(M,S_{M},(M_{I},t_{1}))\rightarrow\exists f^{2}_{process}(M,S_{M},(M_{I},t_{1}))$
$(Eq^{2}(S_{M_{0}}(M_{0},t_{2}),f^{2}_{process}(M,S_{M},(M_{I},t_{1}))))\wedge Output^{3}(M,t_{2},S_{M_{0}}(M_{0},t_{2}))$
indicates that, at time $t_{1}$, if $S_{M_{I}}\left(M_{I},t_{1}\right)$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate ${Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$, then $M_{D}$ can process $S_{M_{I}}\left(M_{I},t_{1}\right)$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$ to yield the output carrier state $S_{M_{O}}\left(M_{O},t_{2}\right)$, which is then output at time $t_{2}$. This demonstrates the model’s ability to process and output information, where $M_{O}$ denotes the output device of $M$.
@@
It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},{To}_{h})$, and ${State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ all represent sets, ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas. Regardless of the situation, both cases satisfy Definition B.1 concerning finite automaton states. Therefore, subsequent discussions will encompass both cases without further elaboration.
@@
where ${\psi o}_{1}\trianglelefteq Io({\varphi o}_{1})$, ${\psi o}_{2}\trianglelefteq Io({\varphi o}_{2})$, and ${\psi o}_{3}\trianglelefteq Io({\varphi o}_{3})$ are the enabling mappings of ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$, respectively. This process establishes $M$ as a typical finite automaton and an operational information system.

### v12

@@ 3.1 Information as Enabling Mapping
Figure 1: Schematic Diagram of Enabling Mapping, Noisy Information, and Information Chain.
Based on Appendix A’s higher-order formal system $\mathcal{L}$ and the result that virtually any state set can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ [33], we obtain a more robust theoretical foundation for defining information:
@@
Definition 2 (Definition of Information) Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\mathbb{\subseteq O}\cup\mathbb{S}$, time of occurrence $T_{h}\mathbb{\subseteq T}$, ontological state set $S_{o}\left(o,T_{h}\right)$, carrier $c\mathbb{\subseteq O}$, reflection time $T_{m}\mathbb{\subseteq T}$, and carrier reflection set $S_{c}\left(c,T_{m}\right)$ are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ or $I=\left\langle o,T_{h},S_{o},c,T_{m},S_{c}\right\rangle$.
The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world [34]. Definition 2 subsumes Shannon’s model: source events and channel codes are states, and source-to-channel is an enabling mapping; OIT thus generalizes classical information theory [35].
@@ 3.2 Noisy Information
[22] states that an information instance $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ admits an inverse mapping $I^{-1}(S_{c}(c,T_{m}))$. In practice, however, information realization is rarely immune to noise or other external factors. Consequently, applying $I^{-1}$ may fail to fully recover $S_{o}(o,T_{h})$; instead, it yields a state corrupted by noise. This motivates the following notion.
Definition 3 (Noisy Information) For information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, the inverse $I^{-1}(S_{c}(c,T_{m}))$ is another state of the ontology $o$ at time $T_{h}$, denoted by $\widetilde{S_{o}}\left(o,T_{h}\right)$, and given by
@@
where $S_{o}^{-}$ is the noise-loss state with $S_{o}^{-}\subseteq S_{o}$, and $S_{o}^{+}$ is the noise-addition state with $S_{o}{\cap S}_{o}^{+}=\varnothing$. The carrier state $S_{c}(c,T_{m})$ can thus be viewed as the image of $\widetilde{S_{o}}\left(o,T_{h}\right)$ under an enabling mapping:
$\widetilde{I}:\widetilde{S_{o}}\left(o,T_{h}\right)\rightrightarrows S_{c}\left(c,T_{m}\right)$ (3.3)
which is the noisy information corresponding to $I$, while $I$ itself is the noise-free information.
Fig. 1(b) depicts the structural characteristics of noisy information. In Shannon’s terms, noise transforms the source state via deletion and addition before transmission; the resulting channel output realizes the noisy carrier state [35]. This structural decomposition will support our TVD bounds by isolating loss and addition masses.
@@ 3.3 Information Chain
[22] uses composition to discuss the transitivity of information. Specifically, if $I_{o}:S_{o}(o,T_{h})\rightrightarrows S_{c_{1}}(c_{1},T_{m_{1}})$ and $I_{1}:S_{c_{1}}(c_{1},T_{m_{1}})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ are two information instances, then $I_{2}:S_{o}(o,T_{h})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ is also an information instance, interpreted as $I_{0}$ being transmitted from carrier $c_{1}$ to carrier $c_{2}$. This yields the notion of an information chain.
Let $\left\{I_{i}:S_{o_{i}}(o_{i},T_{h_{i}})\rightrightarrows S_{c_{i}}(c_{i},T_{m_{i}})\middle|i=1,\cdots,n\right\}$ be a sequence of information mappings, where for any $i>1$, it holds that $o_{i}=c_{i-1}$, $T_{h_{i}}=T_{m_{i-1}}$, and $S_{o_{i}}\left(o_{i},T_{h_{i}}\right)=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$. In this case, we say that these mappings constitute an information chain. Moreover, for any 1 $\leq i\leq j\leq n$, $I_{ij}:S_{o_{i}}(o_{i},T_{h_{i}})\rightrightarrows S_{c_{j}}(c_{j},T_{m_{j}})$ also qualifies as information according to Definition 2.
Transmission constitutes a fundamental mode of information flow and application. In many cases, information undergoes more complex flows through chain-like transmission processes. Fig. 1(c) intuitively illustrates such a process. For any pair of states within this process, the latter can be regarded as the result of the enabling mapping from the former, and this enabling mapping can be used to reasonably explain the causation of the latter state. This provides a theoretical foundation for the interpretability of the model.
3.4 Finite Automata as Formal State Systems
Any practical computing or ML system can be formalized as a finite automaton with state, input, output, and transition functions over discrete times (Appendix B). We represent its behavior by a state set of WFFs that captures inputs, outputs, and transitions, together with predicates and functions for assignability, learnability, and processability. This provides the operational substrate for our lifecycle formalization in Section 4 and for the logical definitions and theorems in Section 5.
@@ 4 Meta-Framework for Machine Learning Theory (MLT-MF)
The core of ML is the artificial intelligence model, which can be regarded as a finite automaton as well as an information system. The processes of operation, learning, testing, and application for such a model all fundamentally depend on information within the system. Therefore, by employing the concepts of information model and finite automaton state, the process of ML and its applications can be unified into five primary steps of information realization and functioning. Fig. 2 presents the chain-structured architecture of causal information progression across the five stages. The subsequent subsections provide formal expressions of the information states for each stage. Undoubtedly, these pieces of information are the fundamental basis for explaining model behavior and outcomes. They transcend specific algorithms or models and provide a new theoretical framework—MLT-MF—for addressing a range of critical issues such as ML interpretability.
Figure 2: Causal Chains of Multiple Information Mappings in Machine Learning.
@@ 4.1 Construction of Initial Information
For the first stage in Fig. 2, suppose the model under study is denoted as $M$, and let $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ be the set of time points at which $M$, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order. According to the perspective of OIT, the creation of $M$ is essentially the realization of information:
$Io:S_{M_{D}}(M_{D},To_{h})\rightrightarrows S_{M}(M,t_{1})$ (4.1)
where the ontology $M_{D}$ of $Io$ is the designer’s conceptualization of $M$, ${To}_{h}$ is the time at which the designer conceives $M$, and $S_{M_{D}}\left(M_{D},{To}_{h}\right)$ is the initial finite automaton state that must be realized at this conceptual stage. The carrier $M$ of $Io$ is the model itself; the reflection time $t_{1}$ is when $M$ is constructed according to $S_{M_{D}}\left(M_{D},{To}_{h}\right)$; and $S_{M}\left(M,t_{1}\right)$ is the formalization of the actual physical state of $M$ upon its completion. Thus, $max({To}_{h})<max(t_{1})$.
@@
$\varphi o_{2}=$ $\mathit{State}^{3}(M_{D},t_{1},\Phi^{2}(M_{D},To_{h}))\land$ (4.3)
$\mathit{Input}^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
$\mathit{Learnable}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{\mathit{learn}}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$
$\left(\mathit{Eq}^{2}(\Phi^{2}(M_{D},t_{2}),f_{\mathit{learn}}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1})))\right)\land$
$\mathit{State}^{3}(M_{D},t_{2},\Phi^{2}(M_{D},To_{h}))$
@@
$\varphi o_{3}=$ $\mathit{State}^{3}(M_{D},t_{1},\Phi^{2}(M_{D},To_{h}))\land$ (4.4)
$\mathit{Input}^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
$\mathit{Processable}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{\mathit{process}}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$
$\left(\mathit{Eq}^{2}(S_{M_{O}}(M_{O},t_{2}),f_{\mathit{process}}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1})))\right)\land$
$\mathit{Output}^{3}(M_{D},t_{2},S_{M_{O}}(M_{O},t_{2}))$
indicates that, at time $t_{1}$, if $S_{M_{I}}\left(M_{I},t_{1}\right)$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate ${Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$, then $M_{D}$ can process $S_{M_{I}}\left(M_{I},t_{1}\right)$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$ to yield the output carrier state $S_{M_{O}}\left(M_{O},t_{2}\right)$, which is then output at time $t_{2}$. This demonstrates the model’s ability to process and output information, where $M_{O}$ denotes the output device of $M$. When output information must not violate relevant ethical guidelines, $f_{process}^{2}$ must implement the corresponding safety assurance functions.
@@
It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},{To}_{h})$, and ${State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ all represent sets, ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$ may be considered as three WFFs, or—when decomposed with respect to the elements of each set—as three sets of more granular WFFs. Regardless of the situation, both cases satisfy Definition B.1 concerning finite automaton states. Therefore, subsequent discussions will encompass both cases without further elaboration.
@@
where ${\psi o}_{1}$, ${\psi o}_{2}$ and ${\psi o}_{3}$ are the enabling mappings of ${\varphi o}_{1}$, ${\varphi o}_{2}$ and ${\varphi o}_{3}$, respectively. This process establishes $M$ as a typical finite automaton and an operational information system.